OpenStudy (anonymous):

Kenneth deposited $750.00 into an account and fifteen years later, the account had earned $787.50 in interest. What simple interest rate did the account earn?

OpenStudy (anonymous):

The account earned 105% from the original value.

OpenStudy (anonymous):

Gained 787.50-750/750%=5% use simple interest formula for calculate the rate SI=PTR/100

OpenStudy (anonymous):

SI=PTR/100 787.50=750*15*r/100 r=7